# Federated Continual Learning: Concepts, Challenges, and Solutions

### v1

@@ Abstract
journal: Computer Science Review
@@ 1 Introduction
Currently, the literature offers only a couple of surveys on FCL. The first published survey categorizes methods based on knowledge fusion mechanisms and introduces synchronous and asynchronous FCL frameworks[6]. In addition, a recent preprint focuses on FCL within Edge-AI environments, with an emphasis on task-specific classifications and hardware-algorithm co-design [12].While these surveys have contributed valuable insights on FCL, focusing on knowledge fusion, task-specific classifications, and Edge-AI contexts. However, they lack a comprehensive taxonomy of key challenges, such as heterogeneity, communication overhead, and model stability, and offer limited cross-domain insights from FL and CL. This survey aims to build on prior works to address these gaps.
This survey addresses the pressing need for a detailed examination of FCL, motivated by its relevance to real-world applications and the open problems in the field. By systematically exploring how FL and CL challenges converge in FCL, this work aims to provide researchers with a clear understanding of the domain’s current state and future directions. Specifically, it highlights critical issues such as catastrophic forgetting in federated settings, non-stationary client environments, and the interplay of privacy, communication, and computational constraints. By bridging the gap between FL and CL, this survey aims to serve as a foundational reference for developing robust and scalable FCL systems, ultimately contributing to the broader adoption of intelligent, adaptive, and decentralized AI solutions.
@@ 2 Federated Continual Learning
The data-driven world requires learning from continuous data in a distributed manner, while respecting privacy. FCL emerges as a robust solution for scenarios involving both collaborative learning and streaming data. FCL seamlessly integrates FL principles with the capability to manage continuous data streams, offering a versatile and efficient approach to contemporary data challenges. By leveraging FCL, it is possible to achieve the dual objectives of collaboratively training models without centralized data collection and maintaining up-to-date models through continual learning from streaming data.
@@ 3.1 Heterogeneity
In collaborative learning, different clients contribute to training, which can enhance overall training performance by leveraging server computational power and improving convergence rates and globalization [21, 22]. However, this diversity can introduce heterogeneity, stemming from factors such as geographical differences, variations in data source distribution, differing data formats, inconsistent data quality, variations in data collection methods, discrepancies in labeling practices, and differences in feature representation [23]. Heterogeneity, which refers to the presence of differences and diversity within a set of elements, is one of the prominent challenges in collaborative learning. [24] highlights the severe impacts of device and behavioral heterogeneity on model performance, showing substantial degradation in model quality and fairness. Another study on dialogue generation focuses on branch homogeneity problems in collaborative learning and proposes methods to enhance branch heterogeneity for improved performance [25]. Tailored strategies are essential for managing heterogeneity, as it can hinder the effectiveness of collaborative learning. Effective management is crucial to fully realize the potential of collaborative learning. This subsection explains different types of heterogeneity and proposes methods to address them effectively, categorizing heterogeneity into two primary types: data heterogeneity and system heterogeneity.
@@ 3.1.1 Statistical Heterogeneity
Differences in input space occur when client $i$ and client $j$ contribute to the training and their local datasets have different distributions $P(x_{i})\neq P(x_{j})$. On the other hand, behavioral differences refer to variations in the way different clients generate, collect, or label their data, leading to different perspectives on events for the machine learning models $P(y_{i}|x_{i})\neq P(y_{j}|x_{j})$. These differences can impact the performance of machine learning models trained on such datasets. Imagine a smart home energy management system deployed across households in different regions, each acting as a client contributing data to a global energy optimization model. Client $i$ in Region A (temperate climate) has energy usage peaks during winter for heating and summer for cooling, with high Heating, Ventilation, and Air Conditioning (HVAC) consumption. Client $j$ in Region B (tropical climate) has consistently high energy usage for cooling year-round. Thus, their energy consumption patterns differ significantly ($P(x_{i})\neq P(x_{j})$). Households in Region A follow consistent daily routines, leading to predictable energy usage peaks in the morning and evening. In contrast, households in Region B have varied routines, with unpredictable energy usage patterns throughout the day. These behavioral differences lead to variations in energy consumption metrics $P(y_{i}|x_{i})\neq P(y_{j}|x_{j})$. To address Non-IID data as a prominent challenge in distributed systems, the proposed methods are reviewed here to clarify the importance of each way.
To address changes in the input space through clients, three different methods are proposed. Firstly, domain transformation involves techniques designed to align the data distributions from multiple clients to create a more unified input space for a machine learning model. One approach focuses on domains with particular features, leveraging the unique characteristics of each domain to enhance performance. For instance, studies [93, 94, 95] address the issue that each domain may have its own unique set of features to characterize samples, leading to incompatibilities across domains, and they develop methods to extract a common feature representation. To talk in detail, deep learning models are utilized for learning complex relations between features [96], and domain generalization [97]. Moreover, [98] proposed an approach in which feature transformation and KL divergence minimization are used to align source and target data distributions. Another approach, performed in [99, 100, 101, 102, 103, 104], is domain factorization, which decomposes data into shared and domain-specific components using methods such as matrix factorization. This isolates common patterns and unique features, allowing models to better understand the underlying structure. These techniques improve the robustness and effectiveness of machine learning models across diverse datasets.
@@ 3.1.2 Conceptual Heterogeneity
When different clients participate in training a model their differences in the components and capabilities lead to system heterogeneity. This heterogeneity can be broadly categorized into two main subsets: hardware heterogeneity and software heterogeneity. Hardware heterogeneity encompasses differences in computational power, memory and storage capacities, and network connectivity among the devices. For example, some devices might have high-end processors capable of fast computations, while others may have limited processing power, leading to synchronization issues during model training. Similarly, variations in memory and storage capacities can result in some nodes struggling with large datasets or complex models, creating bottlenecks. Diverse network connectivity also plays a role, with some devices connected via high-speed broadband and others relying on slower mobile data, affecting data transmission rates and overall system efficiency. On the other hand, software heterogeneity involves the diversity of operating systems, software versions, and programming languages or frameworks used across different nodes. Devices in the system might run on various operating systems such as Android, iOS, Windows, or Linux, requiring robust software solutions that are compatible across all platforms. Different versions of the same software or libraries can cause execution inconsistencies, necessitating careful version control and updates. Additionally, the use of various programming languages and frameworks, such as Python, Java, TensorFlow, or PyTorch, requires interoperability solutions to integrate these diverse technologies seamlessly. Addressing both hardware and software heterogeneity is crucial for the efficient and effective operation of FL and distributed systems, ensuring consistent performance and reliability across all nodes.
@@ 3.4 Model Stability
When multiple clients participate in the training process, it is crucial to consider the global model’s ability to maintain consistent and reliable performance as it is updated over multiple rounds with data from various distributed sources. This is the essence of model stability, which involves managing the inherent variations and challenges posed by the FL environment. Key aspects of model stability include consistent performance, convergence, and scalability. Consistent performance means that the model should show stable performance metrics across different training rounds, avoiding significant fluctuations. Convergence ensures that the model parameters and performance metrics gradually stabilize, indicating effective learning and adaptation. Scalability involves the model’s ability to maintain stability and performance as the number of participating clients increases, handling the added complexity and variability without degradation. Together, these factors contribute to a robust and reliable global model that generalizes well across diverse and distributed client data.
Several techniques can be employed to enhance model stability in FL, including Federated Averaging (FedAvg), regularization, learning rate scheduling, data augmentation, client selection strategies, and adaptive aggregation. The FedAvg algorithm, introduced by [3], enhances model stability by allowing clients to perform local updates on their data and then averaging these updates to form a global model. This method significantly reduces communication overhead and maintains robust performance, even with non-IID data distributions, contributing to consistent and reliable model training across multiple rounds.
Applying regularization techniques (e.g., L2 regularization, dropout) is another approach to prevent overfitting and ensure smoother convergence. [160] demonstrates that dynamic regularization can significantly improve the robustness and effectiveness of FL models. Learning rate scheduling, which involves dynamically adjusting the learning rate, helps ensure steady progress in model updates and prevents overshooting or slow convergence.
To address model stability, enhancing the diversity of training data on clients through data augmentation is essential. [161] proposes a FL method based on data augmentation, abbreviated as FedM-UNE, which implements the classic data augmentation technique MixUp in federated scenarios without transferring raw data. This approach aims to enhance model stability and performance by leveraging the benefits of data augmentation, while maintaining data privacy and minimizing communication overhead.
Moreover, selecting a representative subset of clients for each training round is important to balance the contributions from different data distributions, which can tackle model instability. [162] proposes a client selection strategy that considers both the computational capabilities and communication bandwidth of clients to optimize the training process. By prioritizing clients with adequate resources, the method aims to reduce training time and improve the overall efficiency and performance of FL.
Using adaptive methods for aggregating model updates, such as weighting updates based on the quality or quantity of client data, is another approach for tackling the model instability problem. [163] proposes an approach that includes adaptive aggregation methods and client selection strategies to optimize the utilization of computational and communication resources. By adapting to the varying resource constraints of edge devices, the proposed framework aims to improve the efficiency, scalability, and performance of FL.
@@ 4.2 Solutions
4.2.1 Datasets
To be able to use these techniques in contributive systems, our data requires some features to havewhich can be grouped into three categories. Most of the FCL research uses different publicly available datasets across various domains to validate any particular learning strategy for their challenges, including concept drift and catastrophic forgetting. Table 3.5 summarizes a comprehensive list of datasets and repositories on the basis of dimensionality, data format, type of drift, and respective repositories. The table also summarizes datasets to observe gradual and abrupt drifts with their dimensional parameters in terms of pixels, channels, or features. By providing this in-depth overview, Table 3.5 acts as a foundational point of reference for the selection of datasets that correspond to certain FCL research objectives.
@@ 5 Challenges of Federated Continual Learning
[33] BGD/SGD /mini-batch SGD DynCentAsyncFedAvg SstrSAGA[208] MS/RC**
@@
SOFed [37] BYOL/ SimCLR/MoCo Dyn CentSyncFedAVGU CS HCFLF
@@
AFAFed [11] IoTNet [210] /RMNv2 Stc CentAsyncEASGD S* CO/H/RC**
@@
TARGET [90]✓ResNetDynCentSyncTargetS*PPLD/CFLF/GF
@@ 5.1.1 Catastrophic Forgetting
Knowledge distillation is a widely used technique for fusing knowledge in FCL [220]. The core idea involves transferring the knowledge from an already trained teacher model to a student model. Essentially, the teacher model guides the student to replicate the same mapping relationship by providing the label $Y$ for the sample $X$, similar to the way humans teach. The continual distillation mechanism uses distillation techniques at multiple scales and employs contrastive learning to transfer knowledge from old tasks to new tasks while preserving the learned knowledge [64]. Federated Not-True Distillation (FedNTD) [221] focuses on preserving the global model’s knowledge by only considering not-true classes during distillation, thus maintaining the global perspective without directly accessing local data. The Enhancer distillation method [84] incorporates a small module, the enhancer, to absorb and communicate new knowledge, thereby mitigating the imbalance between old and new knowledge. Lastly, exemplar-free distillation [81] avoids the use of a rehearsal buffer by employing distillation methods that do not require storing past data samples.

### v2

@@ Abstract
journal: Neurocomputing
@@ 1 Introduction
Table 1 provides a comprehensive overview of recent studies in FCL, FL, and CL, focusing on how they address key challenges and propose corresponding solutions. For instance, [6] categorizes FCL methods based on knowledge fusion mechanisms and centers on the issue of forgetting in both synchronous and asynchronous FCL settings. Similarly, [12] investigates FCL from the perspective of three key task characteristics, including federated class continual learning, federated domain continual learning, and federated task continual learning. Other works such as [2, 9, 13] partially address FCL by exploring topics like online FL, CL within FL. For instance, [2] examines the overlap between federated, transfer, and online learning, aligning with FCL principles. [9] addresses data heterogeneity in FL, highlighting challenges from non-IID data and concept drift, and links these issues to CL. Similarly, [13] investigates catastrophic forgetting across domains, including its impact within FCL settings. While these surveys have contributed valuable insights on FCL, focusing on knowledge fusion, task-specific classifications, and Edge-AI contexts. However, they lack a comprehensive taxonomy of key challenges, such as heterogeneity, communication overhead, and model stability, and offer limited cross-domain insights from FL and CL.
As shown in the table, all recent CL surveys discuss catastrophic forgetting, a core challenge in CL. However, [14] and [15] also provide an in-depth analysis of concept drift, which is less commonly addressed. Furthermore, [1] and [5] highlight dilemmas in CL through statistical analyses, visualizations, and diagrams, offering valuable insights into practical limitations.
Privacy is a fundamental concern in FL, and all recent FL surveys acknowledge it. Notably, [4, 16] deeply examine various attack vectors and defense mechanisms in FL systems. Heterogeneity is widely recognized as a fundamental challenge in FL. For instance, [17] specifically addresses data heterogeneity by examining issues such as class imbalance and missing classes. In a broader context, [18] highlights heterogeneity as a critical obstacle in FL, analyzing its impact across data, model, and system levels. Additionally, [19] offers an extensive overview of heterogeneity in FL, covering a wide range of factors including data, model, task, communication, and device heterogeneity.
Finally, while model stability, communication overhead, and resource constraints are recognized as major challenges in FL, they remain underexplored in the literature.
This survey provides a structured analysis of FCL by examining the intersection of challenges from FL and CL. It identifies key challenges, reviews existing methodologies, and explores future research directions. The main contributions are as follows:
  • We systematically examine how the challenges of FL and CL contribute to FCL, providing researchers with a comprehensive understanding of the current state of the domain and identifying key future research directions.
  • We identify and categorize the fundamental challenges in FCL, distinguishing between global and local issues, and trace their origins to specific challenges inherent in FL and CL. This approach enhances the clarity of the underlying factors contributing to FCL challenges, helping researchers and practitioners to understand how these challenges manifest and interact at both global and local levels. By uncovering the root causes, our analysis paves the way for more targeted solutions and effective strategies, ultimately improving the performance, scalability, and robustness of FCL systems.
  • We review and analyze current studies in FCL, examining the challenges addressed in each work. Based on this, we categorize and evaluate the existing solutions to FCL challenges, offering a clearer understanding of how different approaches that tackle these issues and identifying areas for future development.
Table 1: Exploring challenges in federated learning, continual learning, and federated continual learning. This table provides an overview of recent studies. Here L and G stand for local catastrophic forgetting, and global catastrophic forgetting, respectively.
Challenges Studies Subcategory FCL CL FL
Ours [12] [6] [2] [9] [13] [14] [20] [1] [21] [15] [5] [22] [16] [23] [19] [24] [25] [18] [26] [17] [4]
Heterogeneity Statistical Heterogeneity
Conceptual Heterogeneity
System Heterogeneity
Resource Constraints
Communication Overhead
Model Stability
Privacy Preservation Homomorphic Encryption
Differential Privacy
Secure Multi-Party Computation
Secure Aggregation
Concept Drift Virtual Drift
Real Drift
Hybrid Drift
Dilemmas
Catastrophic Forgetting Local Forgetting
Global Forgetting
Global Model Overfitting
Negative Knowledge Transfer
Significant Accuracy Loss
Updating Frequency
@@ 2 Federated Continual Learning
The data-driven world requires learning from continuous data in a distributed manner, while respecting privacy. FCL emerges as a robust solution for scenarios involving both collaborative learning and streaming data. FCL seamlessly integrates FL principles with the capability to manage continuous data streams, offering a versatile and efficient approach to contemporary data challenges. By leveraging FCL, it is possible to achieve the dual objectives of collaboratively training models without centralized data collection and maintaining up-to-date models through CL from streaming data.
@@ 3.1 Heterogeneity
In collaborative learning, different clients contribute to training, which can enhance overall training performance by leveraging server computational power and improving convergence rates and globalization [35, 36]. However, this diversity can introduce heterogeneity, stemming from factors such as geographical differences, variations in data source distribution, differing data formats, inconsistent data quality, variations in data collection methods, discrepancies in labeling practices, and differences in feature representation [37]. Heterogeneity, which refers to the presence of differences and diversity within a set of elements, is one of the prominent challenges in collaborative learning. [38] highlights the severe impacts of device and behavioral heterogeneity on model performance, showing substantial degradation in model quality and fairness. This subsection explains different types of heterogeneity and proposes methods to address them effectively, categorizing heterogeneity into two primary types: data heterogeneity and system heterogeneity.
@@ 3.1.1 Statistical Heterogeneity
Differences in input space occur when client $i$ and client $j$ contribute to the training and their local datasets have different distributions $P(x_{i})\neq P(x_{j})$. On the other hand, behavioral differences refer to variations in the way different clients generate, collect, or label their data, leading to different perspectives on events for the machine learning models $P(y_{i}|x_{i})\neq P(y_{j}|x_{j})$. These differences can impact the performance of machine learning models trained on such datasets. To address non-IID data as a prominent challenge in distributed systems, the proposed methods are reviewed here to clarify the importance of each approach.
To address changes in the input space through clients, three different methods are proposed. Firstly, domain transformation involves techniques designed to align the data distributions from multiple clients to create a more unified input space for a machine learning model. One approach focuses on domains with particular features, leveraging the unique characteristics of each domain to enhance performance. For instance, studies [106, 107, 108] address the issue that each domain may have its own unique set of features to characterize samples, leading to incompatibilities across domains, and they develop methods to extract a common feature representation. To talk in detail, deep learning models are utilized for learning complex relations between features [109], and domain generalization [110]. Moreover, [111] proposed an approach in which feature transformation and KL divergence minimization are used to align source and target data distributions. Another approach, performed in [112, 113, 114, 115, 116], is domain factorization, which decomposes data into shared and domain-specific components using methods such as matrix factorization. This isolates common patterns and unique features, allowing models to better understand the underlying structure. These techniques improve the robustness and effectiveness of machine learning models across diverse datasets.
@@ 3.1.2 Conceptual Heterogeneity
3.1.3 System Heterogeneity
System heterogeneity in FL arises from differences in client hardware and software, impacting training efficiency and reliability. Hardware heterogeneity includes variations in computational power, memory, storage, and network connectivity. High-end devices process data efficiently, while resource-constrained nodes may struggle with large models, causing synchronization delays. Network disparities further affect data transmission rates and system performance.
Software heterogeneity stems from diverse operating systems (e.g., Android, iOS, Windows, and Linux), software versions, and programming frameworks (e.g., TensorFlow, PyTorch, and Java). Inconsistent software versions can lead to execution issues, requiring strict version control and interoperability solutions. Addressing these challenges is essential for ensuring seamless operation, consistent performance, and robustness in FL and distributed systems.
@@ 3.4 Model Stability
When multiple clients participate in the training process, it is crucial to consider the global model’s ability to maintain consistent and reliable performance as it is updated over multiple rounds with data from various distributed sources. Key aspects include consistent performance (avoiding large fluctuations), convergence (gradual stabilization of parameters), and scalability (handling an increasing number of clients without performance degradation). These factors contribute to a robust model that generalizes properly across diverse client data.
Several techniques enhance model stability in FL, including Federated Averaging (FedAvg), regularization, learning rate scheduling, data augmentation, client selection, and adaptive aggregation. FedAvg [3] improves stability by averaging local updates, reducing communication overhead, and handling non-IID data. Regularization techniques (e.g., L2 regularization, and dropout) help to prevent overfitting and support smoother convergence, with dynamic regularization shown to enhance FL robustness [172]. Learning rate scheduling ensures steady progress and prevents instability.
Data augmentation enhances training diversity, while preserving privacy. For instance, FedM-UNE [173] applies augmentation without transferring raw data, improving model stability. Client selection strategies balance contributions from different data distributions by prioritizing clients with sufficient computational and communication resources [174], optimizing training efficiency. Adaptive aggregation methods, which weight model updates based on data quality or quantity, further improve robustness and scalability [175]. These strategies collectively contribute to a more stable and reliable FL framework.
@@ 4.2 Solutions
4.3 Catastrophic Forgetting and Continual Learning Dilemmas
Catastrophic forgetting refers to the tendency of a model to forget previously learned information when acquiring new knowledge. Figure3 highlights catastrophic forgetting in CL by visualizing task-wise accuracy evolution in CL using Elastic Weight Consolidation (EWC) algorithm on CIFAR-10 dataset.
To effectively assess catastrophic forgetting, researchers employ several well-established metrics that quantify the extent to which a model loses knowledge of previously learned tasks, when acquiring new ones. Figure 4 illustrates three key metrics commonly used in CL literature.
First, the Forgetting Measure (FM) is computed individually for each task and reflects the drop in performance from the task’s highest accuracy (immediately after training) to its final accuracy (after learning all tasks). A higher FM value indicates a greater degree of forgetting during training [218].
Second, Average Forgetting (AF) aggregates the forgetting measures across all tasks, providing a single scalar value that summarizes the model’s forgetting behavior. This global metric is particularly useful for comparing the effectiveness of different CL methods [218].
Lastly, Backward Transfer (BWT) captures how learning new tasks influences the performance on previous tasks. BWT can be either positive, indicating knowledge transfer that improves earlier task performance, or negative, signifying catastrophic forgetting [218, 219, 220].
The broader implications of catastrophic forgetting are further explored in Section 5, which categorizes challenges and solutions within the FCL domain.
In contrast to catastrophic forgetting, which reflects a model’s loss of previously acquired knowledge, intransigence denotes the model’s difficulty in learning new tasks. This typically results from an excessive bias toward earlier tasks, often introduced by strong regularization techniques or heavy reliance on prior task-specific knowledge [218]. Such constraints promote stability but suppress plasticity, limiting the model’s capacity to adapt to novel information. The coexistence of forgetting and intransigence underscores the fundamental stability–plasticity dilemma in CL [1, 5, 218, 221, 222]. This dilemma encapsulates the inherent trade-off between stability, which safeguards learned knowledge, and plasticity, which facilitates adaptation to new experiences (see figure 5).
Figure 3: Task-wise accuracy evolution in CL using EWC on CIFAR-10: The plot illustrates the performance trajectory of a CL model trained using EWC on the CIFAR-10 dataset. The dataset is partitioned into five sequential tasks, each comprising two classes. As training progresses through tasks, the accuracy on earlier tasks progressively declines, highlighting the catastrophic forgetting phenomenon. EWC attempts to mitigate this by preserving important parameters from previous tasks, yet the model still experiences notable performance drops, as indicated by the percentage reductions annotated on the plot. This visualization underscores the challenges of maintaining stability in CL settings, even with regularization-based approaches like EWC.
Figure 4: Measures commonly used for catastrophic forgetting
Figure 5: The Stability–Plasticity Trade-off in Continual Learning: A Dilemma-Centric Perspective
Figure 5 also presents a detailed taxonomy of key dilemmas that arise from the stability and plasticity trade-off in CL. Each row in the table contrasts the goals and implications of plasticity-oriented and stability-oriented approaches across various dimensions. For instance, while plasticity favors quick adaptation and knowledge transfer across tasks, stability ensures retention, privacy preservation, and task-specific accuracy. Dilemmas such as replay versus privacy, transfer versus interference, and task-aware versus task-agnostic learning reflect practical tensions that researchers must address when designing CL systems. Understanding these opposing forces is critical for developing models that can learn continuously over time without succumbing to either forgetting or intransigence.
@@ 5 Challenges of Federated Continual Learning
[46] ✗ BGD/SGD/mini-batch SGD DynCentAsyncFedAvg SstrSAGA[225] MS/RC**
@@
SOFed [50] ✗BYOL/SimCLR/MoCoDyn CentSyncFedAvgU CS HCFLF
@@
AFAFed [11] ✓IoTNet [227]/RMNv2Stc CentAsyncEASGD S* CO/H/RC**
@@
TARGET [103]✓ResNetDynCentSyncTargetS*PPLD/CFGF/LF
@@
[237]✓CNNStcCentSyncFedCILS*H/MSCI/CFGF/LF
FedET[97]✓ResNet/ViT/Bert-Base-Uncased[238]DynCentSync Weighted Averaging(distillation)S*HCFGF/LF
FBL[239]✓Deeplab-v3[240] with ResNetDynCentSync*S*HCI/CFGF/LF
PP-FCIL[241]✓ResNetDynCentSyncFedAvgS*PPCI/CFLF
[242]**DynCentSyncFedAvgSS*MS/H/PPLD/CFGF/GMO/AL
FedMAC [243] MobileNetV2[244] /MobileBERT[245] DynCentSync FedAvg/FedProx /FedRS[246]/FedOpt[247] S*H**
EvoFedIDS[248]✓CNNDynCentSyncFedAvgS*H/RCCFGF/LF
[249]✓ResNetStcCentSyncFedAvgS*H/PPCFLF
FedMGP[250]✓ResNet*CentSync/AsyncFedMGPS*H/PPCFGF/LF
FCLLM-DT[251]✓CNNStcCentSyncFedAvgSSWH/PPLD*
FCL4DD[252]✓ResNetStcCentSyncFedAvgS*HCFGF/LF
[253]✓ResNetStcDcent✗✗S*CO/PPCI/CFAL/LF
MeCo[254]✓ResNet*CentSyncFedAvgS*CO/MS/H/RCCFGF/LF
Loci[255]✗VariedDynCent/DcentSyncFedAvgS*CO/H*NT
SacFL[256]✓LeNet/ResNet/TextCNNDynCentSyncFedAvgS*H/RCCD*
@@ 5.1.1 Catastrophic Forgetting
Knowledge distillation is a widely used technique for fusing knowledge in FCL [257, 97]. The core idea involves transferring the knowledge from an already trained teacher model to a student model. Essentially, the teacher model guides the student to replicate the same mapping relationship by providing the label $Y$ for the sample $X$, similar to the way humans teach. The continual distillation mechanism uses distillation techniques at multiple scales and employs contrastive learning to transfer knowledge from old tasks to new tasks while preserving the learned knowledge [77]. Federated Not-True Distillation (FedNTD) [258] focuses on preserving the global model’s knowledge by only considering not-true classes during distillation, thus maintaining the global perspective without directly accessing local data. The Enhancer distillation method [97] incorporates a small module, the enhancer, to absorb and communicate new knowledge, thereby mitigating the imbalance between old and new knowledge. Lastly, exemplar-free distillation [94] avoids the use of a rehearsal buffer by employing distillation methods that do not require storing past data samples. In the context of a decentralized federated learning framework for joint class and domain continual learning, [253] employs response-based knowledge distillation by aligning the logit vectors of local models with those of an incrementally updated global model. This approach mitigates catastrophic forgetting without requiring access to data from previous nodes.